Use the Integral Test to determine whether the series is convergent or divergent. ∞ n n2 + 2 n = 1 Evaluate the following integral. ∞ 1 x x2 + 2 dx
Answer #1

Consider the following integral 124 ах The objective is to determine whether the series is convergent or divergent using inte

The given function as, The function f (x) co Find the function f is decreasing Differentiate the function f (x) with respect

Find integral t du = lim | | - Use substitution method, u =x2 + 2 du = 2xdx - хах = lim I-ln u =lim |-lnt--ln 1 So, the impro